A project that provides annual cash flows of $21,000 for 5 years costs $67,000 today.

  

a. If the required return is 15 percent, what is the NPV for this project?

  

b. Determine the IRR for this project.
Answer #1

D8 fax =SUM(D2:07) 1 year Cash flows pv@15% Present value 0 $ (67,000.00) 1.0000 $ (67,000.00) 1 $ 21,000.00 0.8696 $ 18,260.

D8 fax =SUM(D2:07) 1 year pv@15% YOU AWN Cash flows -67000 21000 21000 21000 21000 21000 =IRR(B2:37) =C2/1.15 =C3/1.15 =C4/1.